Solve 125x^3 + 343 = 0. Find all complex roots.
How do I do this?

Answer by Jk22(389)

(5x)^3 = -49*7 = -(350 - 7) = -343 = -7^3

x = -7/5 3root(1) = -7/5 3root(exp(2pi*i*m)) = -7/5 exp(2/3*pi*m), m=0,1,2

= -7/5, -7/5exp(2*pi*i/3), -7/5exp(4*pi*i/3)